# An Atomic Skill Library Construction Method for Data-Efficient Embodied Manipulation

### v1

@@ Abstract
Embodied manipulation is a fundamental ability in the realm of embodied artificial intelligence. Although current embodied manipulation models show certain generalizations in specific settings, they struggle in new environments and tasks due to the complexity and diversity of real-world scenarios. The traditional end-to-end data collection and training manner leads to significant data demands, which we call “data explosion”. To address the issue, we introduce a three-wheeled data-driven method to build an atomic skill library. We divide tasks into subtasks using the Vision-Language Planning (VLP). Then, atomic skill definitions are formed by abstracting the subtasks. Finally, an atomic skill library is constructed via data collection and Vision-Language-Action (VLA) fine-tuning. As the atomic skill library expands dynamically with the three-wheel update strategy, the range of tasks it can cover grows naturally. In this way, our method shifts focus from end-to-end tasks to atomic skills, significantly reducing data costs while maintaining high performance and enabling efficient adaptation to new tasks. Extensive experiments in real-world settings demonstrate the effectiveness and efficiency of our approach.
@@ 1 Introduction
Embodied intelligence, primarily referring to “embodied artificial intelligence”, has seen significant advancements in the era of generative AI. By mapping multimodal data such as text, images, and speech into a unified semantic continuous vector space, this achievement enables domain-independent cross-modal integration. This space closely links the semantic discrete symbol space with the feature continuous vector space, providing a new opportunity for embodied intelligence to develop into a general form. End-to-end embodied manipulation, especially Vision-Language-Action (VLA) models, have continuously achieved significant results due to the availability of embodied data O’Neill et al. (2023) Khazatsky et al. (2024) and the advancement of multimodal technologies, demonstrating increasing generality and generalization, greatly enhancing the potential for service robots to be practically applied.
@@
Although current embodied manipulation models demonstrate some level of generalization in specific settings, they face challenges in adapting to new environments and tasks Black et al. (2024) Li et al. (2024). Notably, current embodied manipulation models are characterized by their end-to-end orientation, where both data collection and model training are designed around specific tasks. The traditional end-to-end approach to data collection and training results in substantial data requirements, a phenomenon we refer to as “data explosion.” In scientific research, a certain number of end-to-end tasks can be manually defined to iteratively improve algorithm performance Team et al. (2024) Liu et al. (2024). However, from the perspective of practical application for general embodied manipulation, the complexity of the physical world and the diversity of real-world tasks severely limit the feasibility of manually defining end-to-end tasks. On the one hand, it is almost impossible to exhaustively define real-world tasks, and each new task requires the collection of corresponding new trajectory data. On the other hand, to achieve satisfying generalization in terms of object, position and scene, a large number of end-to-end trajectory data is required for every task Chi et al. (2023) Kim et al. (2024). Although using synthetic data Mu et al. (2021) Gu et al. (2023) Mu et al. (2024) can reduce the cost of data collection, VLA models trained on synthetic data still face significant challenges when applied in real-world settings. Therefore, as the complexity of tasks increases, the demand for trajectory data also rises, leading to the risk of “data explosion”, which significantly hinders the practical application of embodied manipulation models.
Naturally, the execution of an end-to-end task can be decoupled into the sequential execution of atomic skills Diab et al. (2020) Mao et al. (2024). These atomic skills are more fine-grained and general, allowing reuse across different scenarios and tasks. Furthermore, by our experiments, we found that mapping end-to-end tasksTeam et al. (2024) Kim et al. (2024) into atomic skills helps reduce the amount of data demand and improve the performance of embodied models. Previous methods Kroemer et al. (2021) attempts to execute embodied manipulation at the skill level were limited by manually pre-defined skills, which failed to achieve generality and generalization in practical application scenarios. With advancements in multimodal technologies, we now possess the technical foundation for automatic skill segmentation and real-time planning, termed Vision Language Planning (VLP). VLP can facilitate automatic segmentation during atomic skill training and provide critical low-frequency control signals during inference to plan and guide high-frequency atomic skill execution, greatly facilitating the effective construction and practical application of the atomic skill library. Moreover, due to the rapid development and good performance of VLA models in the field of embodied manipulation, they hold significant potential for constructing atomic skill libraries. This paper proposes a method for constructing atomic skill libraries by combining VLP and VLA, which can significantly reduce data collection costs while maintaining generality and generalization in practical application scenarios. Since VLA, VLP, and atomic skills have the ability to self-update, the atomic skill library can continuously iterate and adapt to a wider range of real-world tasks.
@@
Specifically, the proposed method consists of three main modules, termed as “three-wheeled” method because each module can update as the development of relevant technology. 1. VLP, which decomposes end-to-end tasks into several subtasks based on task and scene understanding and spatial perception intelligence; 2. a high-level semantic abstraction module, responsible for abstracting specific subtasks into a set of general atomic skill definitions; 3. VLA, which is trained on data oriented towards atomic skills to acquire the corresponding atomic skills. This method leverages VLP’s planning and high-level semantic abstraction to automatically define the atomic skill set in a data-driven manner, followed by data collection and VLA fine-tuning to achieve atomic skills. Compared to traditional end-to-end strategies, our method achieves the exponential improvement of data collection efficiency and realizes the cross-task and cross-scenario generality. The contributions of this paper are as follows:
@@
  • Based on this framework, we specifically implement a VLP agent capable of effective task decomposition and real-time planning, along with a semantic abstraction strategy that maps subtasks to a set of general atomic skill definitions, integrated with fine-tuning VLA to realize a practical method for atomic skill library construction.
  • To the best of our knowledge, we are among the first to attempt to address the data explosion in practical applications of embodied end-to-end manipulation method through an atomic skill library approach. Extensive experiments in real-world settings show that our atomic skill library method can significantly reduce data costs while demonstrating excellent task execution capabilities.
@@ 2 Related work
Although VLA models have made significant progress in robotic control, current embodied manipulation approaches still face critical challenges. First, they often require massive amounts of data to learn and generalize complex tasks effectively. Second, they lack mechanisms for failure feedback, which limits their ability to improve and adapt over time. Despite these challenges, our work explores the decomposition of complex tasks into smaller atomic tasks, enabling robots to learn the corresponding atomic skills for each.
@@ 3 Method
Figure 2: The VLP model’s reasoning chain framework based on spatial intelligence information.
@@ 3.2 VLP Agent Wheel
As shown in Fig. 2, given a text-based task instruction and image-based current observation, an off-shelf VLM is introduced to generate a scene description corresponding to the observation image. Considering the spatial complexity of the 3D world, we additionally devise a strategy for spatial relationship awareness. We first use Dino-X Ren et al. (2024), a remarkable object detection model, to detect task-related objects in the observation, outputting the position of each object in the form of bounding boxes. To pursue more precise localization of objects, SAM-2 Ravi et al. (2024) is used to get the meticulous segmentation mask of each object. Then, a rule-based algorithm is carried out to judge the spatial relation between objects. Combining visual perception with spatial intelligence, we put them together with the task instruction into a powerful LLM Team et al. (2023) and prompt it to decompose the task into subtasks and manage their execution order. The speically designed prompt ask the LLM to sequentially perform the following steps:
  • Generate a more specific and detailed task description based on the given visual description, spatial relationships, and original task instructions.
  • Generate a complete execution plan based on the detailed task description.
  • Break down the execution plan into multiple subtasks.
  • Determine the current subtask stage.
  • Specify the subtask that needs to be executed next.
@@ 4 Experiment
1. When collecting trajectory data under the same physical setting, can our method achieve the camparable performance of the end-to-end method with less data?
2. If collecting the same amount of trajectory data, can our method achieve better performance than the end-to-end method?
3. When faced with a new task, can our method still work well?
@@ 4.1 Experimental Setup
We selected four challenging tasks: picking up a bottle and pouring water, taking a banana and placing it on a plate, picking up and placing the pen into the pen holder, and picking up blocks in specified order. We conducted detailed experiments on these four tasks to answer the three questions mentioned above. Figure 3 provides detailed task definitions and visualizations.
@@
For each task, we collect a certain amount of trajectory data to fine-tune the backbone VLA model. Now, we list the number of demonstrations related to each task in the fine-tuning dataset:
For the banana-grasping task, an initial collection of 24 end-to-end trajectory data samples was obtained from 4 different banana positions and 2 different plate positions. To maintain the same data distribution as the end-to-end approach, 12 trajectories for picking up bananas and 6 for placing them were collected. Subsequently, to ensure the same number of trajectories as the end-to-end method, 24 trajectories were gathered for picking up bananas from 8 banana positions, and 24 for placing bananas at 3 plate positions.
For the task of picking up a bottle and pouring water, a total of 27 end-to-end trajectory data samples were initially collected across 3 different bottle positions and 3 different cup positions. To align with the end-to-end method in terms of data distribution, 9 trajectories for picking up bottles and 9 for pouring water were collected. Then, to maintain the same number of trajectories as the end-to-end method, 27 bottle-picking trajectories were collected from 9 different bottle positions, and 27 water-pouring trajectories from 9 different cup positions.
For the task of sequentially picking up blocks, the task was defined as picking up the red, green, and blue blocks in order, ensuring that each block’s placement remains unchanged. Initially, a total of 10 end-to-end trajectory data samples were collected at pre-defined positions. To maintain the same number of trajectories as the end-to-end approach, 10 trajectories were collected for each of picking up the red block, the green block, and the blue block at the same block positions, totaling 30 trajectories.
For the task of picking up the pen and placing it into the pen holder, 24 trajectory samples were initially collected using the end-to-end method across 4 different pen positions and 2 different pen holder positions. To maintain the same data distribution as the end-to-end method, 12 trajectories for picking up the pen and 6 for placing the pen into the holder were initially collected. Then, to match the same number of trajectories as the end-to-end method, 24 trajectories were collected for picking up the pen from 8 different pen positions and 24 for placing the pen into the holder from 3 different pen holder positions.
@@ 4.2 Results Analysis
Q1: From Table 1, it can be observed that our method integrated with Octo or RDT-1B shows comparable performance to corresponding end-to-end methods. For instance, in the task of picking up a bottle and pouring water, when both the bottle and the cup are positioned outside the training set, our method achieves a 20% improvement in success rate. This demonstrates that our approach requires less data under the same distribution while achieving comparable or even better performance, effectively addressing “data explosion.”
Q2: Similarly, Table 1 shows that when fine-tuned on the same amount of demonstrations, our method significantly improves the success rate regardless of Octo or RDT-1B VLA. For example, in the task of placing a banana onto a plate, when both the banana and the plate are positioned outside the training set, our method increases the success rate by 40%. This improvement stems from gathering data from more diverse locations while maintaining the same data volume, thereby enhancing the model’s generalization capabilities.
@@ 5 Conclusion
In this work, we propose a data-driven framework for atomic skill library construction, termed the “Three-wheeled Self-Driven Atomic Skill Library Construction Method”, to address the “data explosion” issue caused by the traditional end-to-end embodied manipulation strategy. Overall, our approach automatically defines and updates a set of atomic skills in a data-driven manner and then realizes these skills through data collection and VLA fine-tuning. Extensive experiments on real-world scenarios and tasks show the data efficiency and generality of our method. We hope this work can inspire future research on significant solutions to “data explosion”.
@@ 6 Acknowledgments
We would like to acknowledge d-robotics.cc and Agilex robotics for their support in providing the experimental sensors and robotic platforms. Additionally, we acknowledge the authors of RDT-1B for their significant contributions to the development of the VLA model. This work was supported by the National Key Research and Development Program of China (NO.2024YFE0211000).

### v3

@@ Abstract
Embodied manipulation is a fundamental ability in the realm of embodied artificial intelligence. Although current embodied manipulation models show certain generalizations in specific settings, they struggle in new environments and tasks due to the complexity and diversity of real-world scenarios. The traditional end-to-end data collection and training manner leads to significant data demands. Decomposing end-to-end tasks into atomic skills helps reduce data requirements and improves the task success rate. However, existing methods are limited by predefined skill sets that cannot be dynamically updated. To address the issue, we introduce a three-wheeled data-driven method to build an atomic skill library. We divide tasks into subtasks using the Vision-Language-Planning (VLP). Then, atomic skill definitions are formed by abstracting the subtasks. Finally, an atomic skill library is constructed via data collection and Vision-Language-Action (VLA) fine-tuning. As the atomic skill library expands dynamically with the three-wheel update strategy, the range of tasks it can cover grows naturally. In this way, our method shifts focus from end-to-end tasks to atomic skills, significantly reducing data costs while maintaining high performance and enabling efficient adaptation to new tasks. Extensive experiments in real-world settings demonstrate the effectiveness and efficiency of our approach.
@@ 1 Introduction
Embodied intelligence, primarily referring to “embodied artificial intelligence”, has seen significant advances in the era of generative AI. By mapping multimodal data, such as text, images, and speech into a unified semantic continuous vector space, this enables domain-independent cross-modal integration. This space closely links the semantic discrete symbol space with the feature continuous vector space, providing a new opportunity for embodied intelligence to develop into a general form. End-to-end embodied manipulation, especially Vision-Language-Action (VLA) models, have shown significant progress due to the availability of embodied data O’Neill et al. (2023) Khazatsky et al. (2024) and the advancement of multimodal technologies, demonstrating increasing generality and generalization, greatly enhancing the potential for service robots to be practically applied.
@@
Although current embodied manipulation models demonstrate a certain generalization in specific settings, they face challenges in adapting to new environments and tasks Black et al. (2024) Li et al. (2024). Notably, current embodied manipulation models are characterized by their end-to-end orientation, where both data collection and model training are performed based on specific end-to-end tasks. In scientific research, a certain number of end-to-end tasks can be manually defined as standards to facilitate the improvement of algorithm performance Team et al. (2024) Liu et al. (2024). However, the diversity and complexity of real-world scenarios make the end-to-end method infeasible for the practical application in general embodied manipulation. On the one hand, real-world tasks are impossible to enumerate, while the end-to-end method fails to extend the learned task abilities to new tasks. On the other hand, as the complexity of a task increases, such as with more complicated procedures, the data requirements of the task must increase to maintain satisfactory performance. These problems make every single task require a significant number of corresponding new data Kim et al. (2024) Liu et al. (2024), leading to the risk of “data explosion”, which significantly hinders the practical application of embodied manipulation models.
Given these challenges associated with end-to-end methods, a natural solution is to break down the execution of an end-to-end task into the sequential execution of atomic skills Diab et al. (2020) Mao et al. (2024). These atomic skills are more fine-grained and general, allowing reuse across different scenarios and tasks. Furthermore, by our experiments, we found that mapping end-to-end tasksTeam et al. (2024) Kim et al. (2024) into atomic skills helps reduce the data requirement and improve the task execution success rate. Although previous work on atomic skills has demonstrated effectiveness, it is limited by the manual predefinition of specific skills. The predefined skill sets are highly limited and incapable of dynamic updating and expansion, thus remaining highly inefficient for general embodied manipulation in the real world. Actually, automatic task decomposition, planning and skill definition are feasible based on today’s powerful unimodal and multimodal understanding and reasoning technologies.
@@
In light of the potential of atomic skills and the limitations of prior arts, our work focuses on developing a practical atomic skill library construction method to achieve data-efficient embodied manipulation in real-world scenarios. Specifically, the proposed method consists of three main modules, termed as “three-wheeled” method because each module can update as the development of relevant technology. Specifically, during training stage, the Vision-Language-Planning (VLP) module, which integrates visual perception, language understanding, and spatial intelligence, is employed to decompose the given task into multiple subtasks. These subtasks are mapped into a set of atomic skill definitions by the high-level semantic abstraction module based on the plasticity and adaptability of the backbone VLA. Corresponding data collection is performed oriented towards the automatically generated atomic skill definitions, which are used for few-shot learning of the VLA module to achieve skill library. Compared to traditional end-to-end strategies, our method achieves the exponential improvement of data collection efficiency and realizes the cross-task and cross-scenario generality. The contributions of this paper are as follows:
@@
  • Based on this framework, we implement a VLP agent capable of effective task decomposition and real-time planning, along with a semantic abstraction strategy that maps subtasks to a set of general atomic skill definitions, integrated with fine-tuning VLA to realize a practical method for atomic skill library construction.
  • To the best of our knowledge, we are among the first to attempt to address the data explosion in practical applications of embodied manipulation methods through an atomic skill library approach. Extensive experiments in real-world settings show that our atomic skill library method can significantly reduce data costs while demonstrating excellent task execution capabilities.
@@ 2 Related work
Although VLA models have made significant progress in robotic control, current embodied manipulation approaches still face critical challenges, particularly the need for massive amounts of data to learn and generalize complex tasks effectively. Despite these challenges, our work explores the decomposition of complex tasks into smaller atomic tasks, enabling robots to learn the corresponding atomic skills for each.
@@ 3.1 Overview
Figure 2: The VLP agent reasoning chain framework based on spatial intelligence information.
@@ 3.2 VLP Agent Wheel
As shown in Fig. 2, given a text-based task instruction and image-based current observation, the Prismatic model Karamcheti et al. (2024), an off-shelf VLM, is introduced to generate a scene description corresponding to the observation image. Considering the spatial complexity of the 3D world, we additionally devise a strategy for spatial relationship awareness. We first use Dino-X Ren et al. (2024), a remarkable object detection model, to detect task-related objects in the observation, outputting the position of each object in the form of bounding boxes. To pursue more precise localization of objects, SAM-2 Ravi et al. (2024) is used to get the meticulous segmentation mask of each object. Then, a rule-based algorithm is carried out to judge the spatial relation between objects. Combining visual perception with spatial intelligence, we put them together with the task instruction into GPT-4 and prompt it to decompose the task into subtasks and manage their execution order. The specially designed prompt asks GPT-4 to sequentially perform the following steps: generate a complete execution plan based on the detailed task description and specify which subtask needs to be executed next.
@@ 4 Experiment
  • When collecting trajectory data under the same physical setting, can our method achieve the comparable performance of the end-to-end method with less data?
  • By collecting the same amount of data, can our method achieve better performance than the end-to-end method?
  • When dealing with new tasks, can our method still work well with few or without additional data?
  • Is our method applicable to different backbone VLA models while maintaining effectiveness and efficiency?
@@ 4.1 Experimental Setup
We selected four challenging tasks to evaluate the performance of our methods from different dimensions, including complex scenarios that the model may encounter in real-world tasks, such as various positions of objects and sophisticated operations. The first three tasks were specifically designed to validate the data efficiency and performance of our method, while the fourth task was used to assess its adaptability to new tasks. Detailed task definitions and visualizations are provided in Figure 3.
@@
We collected trajectory data samples for fine-tuning the backbone VLA model. The number of demos for each task is as follows:
  • Pour Water: 3 different bottle positions and 3 different mug positions, 3 demos for each setting, leading to 27 demos for the end-to-end fine-tuning, 9 and 9 demos for the skill-based VLA fine-tuning.
  • Pick & Place Banana: 4 different banana positions and 2 different plate positions, 3 demos for each setting, leading to 24 demos for end-to-end VLA fine-tuning, 9 and 6 demos for skill-based VLA fine-tuning.
  • Pick & Place Pen: 3 different bottle positions and 3 different mug positions, 3 demos for each different settings, leading to 27 demos for the end-to-end VLA fine-tuning, 9 and 9 demos for the skill-based VLA fine-tuning.
  • Move blocks in order: 10 demos of the Red-Green-Blue order for the end-to-end VLA fine-tuning, 10 demos of moving red, green, and blue block respectively for the skill-based VLA fine-tuning.
@@ 4.2 Results Analysis
Q1: From Table 1, it can be observed that our method integrated with Octo or RDT-1B shows comparable performance to corresponding end-to-end methods. All the success rates of our method are no less than the end-to-end method. For instance, in the task of picking up a bottle and pouring water, when both the bottle and the cup are positioned outside the training set, our method achieves a 20% improvement in success rate. This demonstrates that our approach requires less data under the same distribution while achieving comparable or even better performance, effectively addressing “data explosion.”
Q2: Similarly, Table 1 shows when fine-tuned on the same amount of demonstrations, our method significantly improves the success rate regardless of Octo or RDT-1B VLA. For example, in the task of placing a banana onto a plate, when both the banana and the plate are positioned outside the training set, our method increases the success rate by 40%. This improvement stems from gathering data from more diverse locations while maintaining the same data volume, thereby enhancing the model’s generalization capabilities.
@@
Q4: Table 1 shows improvement of data efficiency and manipulation performance of our method regardless of Octo or RDT-1B compared to their corresponding end-to-end method. Table 2 shows our method can better adapt to new tasks regardless of the backbone VLA. These results indicate that our method can be effectively applied to various VLA models, improving data efficiency, manipulation performance, and adaptability to new tasks.
@@ 5 Conclusion
In this work, we propose a data-driven framework for atomic skill library construction, termed the “Three-wheeled Self-Driven Atomic Skill Library Construction Method”, to address the “data explosion” issue caused by the traditional end-to-end embodied manipulation strategy. Overall, our approach automatically defines and updates a set of atomic skills in a data-driven manner and then realizes these skills through data collection and VLA fine-tuning. Extensive experiments on real-world scenarios show the data efficiency and generality of our method. We hope this work can inspire future research on significant solutions to “data explosion”.
@@ 6 Acknowledgments
We would like to acknowledge d-robotics.cc, Realman robotics and Agilex robotics for their support in providing the experimental sensors and robotic platforms. Additionally, we acknowledge the authors of RDT-1B for their significant contributions to the development of the VLA model. This work was supported by the National Key Research and Development Program of China (NO.2024YFE0211000).